# Connecting the Dots: Evaluating Abstract Reasoning Capabilities of LLMs Using the New York Times Connections Word Game

### v5

@@ Abstract
The New York Times Connections game has emerged as a popular and challenging pursuit for word puzzle enthusiasts. We collect 200 Connections games to evaluate the performance of state-of-the-art large language models (LLMs) against expert and novice human players. Our results show that even the best-performing LLM, GPT-4o, which has otherwise shown impressive reasoning abilities on a wide variety of benchmarks, can only fully solve 8% of the games. Compared to GPT-4o, novice and expert players perform better, with expert human players significantly outperforming GPT-4o. We create a taxonomy of the knowledge types required to successfully cluster and categorize words in the Connections game, revealing that LLMs struggle with associative, encyclopedic, and linguistic knowledge. Our findings establish the New York Times Connections game as a challenging benchmark for evaluating abstract reasoning capabilities in humans and AI systems.
@@
Prisha Samadarshi2 Mariam Mustafa2 Anushka Kulkarni2 Raven Rothkopf2 Tuhin Chakrabarty1††thanks: *Equal contribution. Smaranda Muresan1,2* 1 Department of Computer Science, Columbia University 2 Barnard College {ps3203, mm5970, ajk2256, rgr2124}@barnard.edu, tuhin.chakr@cs.columbia.edu, smara@cs.columbia.edu
@@ 1 Introduction
Word puzzle enthusiasts have become captivated by Connections, an engaging game launched by the New York Times (NYT) in June 2023. This daily game presents players with a 4x4 grid containing 16 words and tasks them to identifying four distinct clusters that link the corresponding four words in each cluster through some shared characteristics (Figure 1 [a] vs [b]). Despite its seemingly straightforward premise, Connections delivers a stimulating linguistic workout that keeps players returning daily to test their mental acuity and semantic savvy. The categories 1 (yellow), 2 (green), 3 (blue), and 4 (purple) are arranged according to ascending levels of difficulty. Category 1 is the most intuitive, while Category 4 is the hardest. For instance, in Figure 1 (b), the most straightforward category is "Conformists" {Followers, Lemmings, Puppets, Sheep}, while the most challenging category includes {Apartment, Insults, Likes, Shovels} and requires the understanding that a single word (in this case, "digs") can have multiple meanings that differ in etymology or sense, depending on the context.
While the task may appear easy, many words clump easily into categories, acting as red herrings. For instance Likes, Followers, Shares, Insult might be categorized as “Social Media Interactions" at first glance. However, the game is designed to promote orthogonal thinking and pushes players to find unusual ways to group things. To group words across proper categories, as shown in Figure 1 (b), a player must reason with various forms of knowledge spanning from Semantic Knowledge (Conformists) to Encyclopedic Knowledge (U.S. cities).
@@
Abstract reasoning represents a person’s ability to solve problems, identify patterns, and work with logical systems Barrett et al. (2018); Johnson et al. (2021); Ji et al. (2022). While the performance of large language models (LLMs) on arithmetic and language-based commonsense reasoning benchmarks has been the subject of recent analyses, it is unclear whether these LLMs possess abstract reasoning capabilities that are often challenging even for humans Xu et al. (2023). Given its nature,
we choose the NYT Connections Game as a test bed for investigating the abstract reasoning capabilities of both humans and large language models (LLMs). We collect 200 distinct Connection games and test the capabilities of four state-of-the-art large language models, namely Google’s Gemini 1.5 Pro Team et al. (2023), Anthropic’s Claude 3 Opus Anthropic (2024), OpenAI’s GPT4-Omni OpenAI (2023), and Meta’s Llama 3 70B AI@Meta (2024) and compare them with human performance.
@@
While all LLMs can partially solve some of the games, their performance is far from ideal. Experimental evidence shows that with few-shot and chain-of-thought prompting, even the best-performing LLM, GPT-4o, can only solve 8% of the games completely. We recruit human players at novice and expert levels of proficiency and compare their performance to GPT-4o. Our results show that the Connections game serves as a challenging benchmark for reasoning, with novice players performing only marginally better than GPT-4o. On the contrary, expert players perform significantly better than GPT-4o in solving games perfectly (Section 5). To better understand the challenging nature of this benchmark, we create a taxonomy of knowledge required to group words into their respective categories (Section 3.2). Our analysis shows that while LLMs are good at reasoning that involves some types of semantic knowledge, they struggle with other types of knowledge such as associative, encyclopedic, or muti-word expressions. Our code and data will be made available to the public at https://github.com/mustafamariam/LLM-Connections-Solver.
@@ 2 Related Work
The growing popularity in Large Language Models have led to an exciting array of research using natural language processing techniques for text-based games. Recent work has studied whether these models can act as players in agentic environment Huang et al. (2024); Wang et al. (2023); Wu et al. (2023); Noever et al. (2020) or conversational setting Qiao et al. (2023). In addition to acting as players researchers have also tested the abilities of transformer based language models in generating games Ammanabrolu and Riedl (2021); Todd et al. (2023); Sudhakaran et al. (2024); Hu et al. (2024); Chen et al. (2023)
@@
Recent research has explored applying large language models (LLMs) and other natural language processing (NLP) techniques to solve and generate text-based puzzles. Wallace et al. (2022) propose automatic ways of solving crossword puzzles by generating answer candidates for each crossword clue using neural question answering models and combining loopy belief propagation with local search to find full puzzle solutions.Zhao and Anderson (2023) use LLMs to tackle and create the weekly Sunday Puzzles featured on NPR. They release PUZZLEQA, a multiple-choice dataset comprising 15 years of on-air puzzles and show that while ChatGPT can solve these questions with an accuracy of around 50% they still struggle at generating novel and engaging puzzles. Unlike Connections game, PUZZLEQA relies on character-level word transformations compared to encyclopedic, associative, or semantic knowledge. Rozner et al. (2021) examined the potential of using "cryptic crossword" clues as an NLP benchmark.
@@
Contemporaneous work from Todd et al. (2024) test OpenAI’s GPT3.5 and GPT4’s abilities in solving the Connection puzzles using Chain-Of-Thought prompting. Our work is similar to theirs in a way that it utilizes the NYT Connections puzzle as a means to investigate the abstract reasoning capabilities of state-of-the-art LLMs. However, in addition to GPT4 we also test 3 other state of the art LLMs (both open and closed weight) as well as humans(both novice and experts). Our contribution not only focuses on the abilities of LLMs to solve the games but we also study the knowledge types required to solve the Connections game. Additionally Todd et al. (2024)’s experimental setup mirrors the way the original Connections game is set up (i.e. one category at a time, with an allotment of 4 incorrect guesses) while we require both human’s and LLMs to provide all categories at once with only one attempt)
Knowledge | Category | Words | |||
Encyclopedic |
[‘CHEERS’, ‘EUPHORIA’, ‘FELICITY’, ‘GLEE’] | |||
JACKS | [‘BLACK’, ‘FROST’, ‘MA’, ‘SPARROW’] | ||||
Semantic | Synonym | COLLEAGUES | [‘ASSOCIATE’, ‘FELLOW’, ‘PARTNER’, ‘PEER’] | ||
Polysemy | WHAT A MOLE CAN BE | [‘ANIMAL’, ‘BIRTHMARK’, ‘SPY’, ‘UNIT’] | |||
Hypernym | PERIOD | [‘AGE’, ‘DAY’, ‘ERA’, ‘TIME’] | |||
Associative | ORIGIN | [‘CRADLE’, ‘FONT’, ‘ROOT’, ‘SOURCE’] | |||
THINGS THAT ARE ORANGE | [‘BASKETBALL’, ‘CARROT’, ‘GOLDFISH’, ‘PUMPKIN’] | ||||
Linguistic | NOUN SUFFIXES | [‘DOM’, ‘ION’, ‘NESS’, ‘SHIP’] | |||
SILENT “W" | [‘ANSWER’, ‘TWO’, ‘WRIST’, ‘WRONG’] | ||||
Multiword Expression | ___WOOD | [‘DOG’, ‘DRIFT’, ‘HOLLY’, ‘SANDAL’] | |||
Combined | CITY HOMOPHONES | [‘DELI’, ‘NIECE’, ‘ROAM’, ‘SOUL’] | |||
SOCIAL MEDIA APP ENDINGS | [‘BOOK’, ‘GRAM’, ‘IN’, ‘TUBE’] |
@@
The word association task Galton (1879) has been used extensively in psychological and linguistic research as a way of measuring connections between words in the mental lexicon. Responses in word association tasks have informed what we know about the structure and organization of semantic memory and the mental lexicon De Deyne and Storms (2008). In this work, we similarly show how one must utilize semantic and associative memories to solve the Connections game.
@@
Chollet (2019) proposed the Abstraction and Reasoning Corpus (ARC), built upon an explicit set of priors designed to be as close as possible to innate human priors and argued that it can be used to measure a human-like form of general fluid intelligence, enabling fair general intelligence comparisons between AI systems and humans. Recently Xu et al. (2023) show that GPT-4 solves only 13/50 of the most straightforward ARC tasks, demonstrating a significant gap in the abstract reasoning capabilities of LLMs. Prior work has also studied abstract reasoning in Neural Networks Barrett et al. (2018) even in the presence of distracting features Zheng et al. (2019). Our work builds upon these and presents the Connections game as a compelling benchmark for abstract reasoning capabilities for LLMs in the presence of distractors.
@@ 3.1 Collection
To gather the necessary data, we found an archival site consisting of all possible answer choices and their corresponding categorizations. As the NYT does not maintain an archive of Connection puzzles, we resorted to an external, third-party site for data collection.111https://tryhardguides.com/nyt-connections-answers/ Our data spans daily problems from the conception of Connections, June 2023, to March 2024. In total, we gather 203 distinct games, out of which 3 are used for few-shot prompting, while the remaining 200 comprise the dedicated test set.
@@ 3.2 Types of Reasoning
Investigating the relationship between words offers insights into both the structure of language and the influence of cognition on linguistic tasks Stella et al. (2018). To solve Connections games, players must draw on certain aspects of word knowledge, such as a word’s meaning. To deepen our understanding, we bucket each category, grouping into the types of knowledge that are primarily required to solve them. Two experts annotate a total of 800 samples coming from 200 games into 6 broader categories. On 8.6% of the 800 samples where they disagree (See examples of disagreement in Appendix B), the experts engaged in discussion Schaekermann et al. (2018); Chen and Zhang (2023); Chen et al. (2019) to arrive at an individual category.
3.2.1 Semantic Knowledge
@@
The majority of instances in the Connections game require possessing knowledge of lexical semantics Cruse (1986), particularly semantic relations such as synonymy (words with the same meaning), hypernymy/hyponymy (relation between a generic terms and its specific instance), and polysemy (many possible meanings for a word). Table 1 shows three examples of groups that use such Semantic Knowledge.
3.2.2 Associative Knowledge
To group words into their respective categories one often needs to think beyond the lexical semantic relations mentioned above. Associative learning Shanks (1995) occurs when an element is taught through association with a separate, seemingly unrelated pre-occurring element. To cluster words using Associative Knowledge, one either needs to focus on the connotative meaning of a word or the shared property that connects several words. For instance, as shown in Table 1, the words Cradle, Root, or Font in their literal sense do not refer to Origin; instead, one needs to rely on their connotative meaning for such a categorization. Similarly, on the surface level, Basketball, Carrot, Goldfish, and Pumpkin are unrelated. However, a shared property that connects them is their orange color.
3.2.3 Encyclopedic Knowledge
@@
We notice that to group certain sets of words into their proper categories, one needs knowledge that spans beyond concepts and relies on entities in the real world found in knowledge bases such as Wikipedia Mihalcea and Csomai (2007). This can be seen in Table 1, where, to bucket the words Black, Frost, Ma, and Sparrow into the category of Jacks, one needs to possess knowledge across various domains: ‘Jack Black’ an American actor, ‘Jack Frost’ a character from English folklore who personifies winter, ‘Jack Ma’ the founder of Alibaba, and ‘Jack Sparrow’ the protagonist of the Pirates of the Caribbean film series. We label this type of knowledge Encyclopedic Knowledge.
3.2.4 Multiword Expressions
Multiword Expressions are complex constructs that interact in various, often untidy ways and represent a broad continuum between non-compositional (or idiomatic) and compositional groups of words Moon (1998). Higher difficulty levels (blue and purple) in the Connections game often require players to recognize that the four words can form a Multiword Expression if combined with an external word. Table 1 shows examples of Multiword Expressions where half of the expressions are given in the form of individual words and the player needs to find the other half to categorize words into the correct group.
3.2.5 Linguistic Knowledge
Linguistic competence Coseriu (1985) is the system of unconscious knowledge that one has when one knows a language. Such competence is often required to classify words into their appropriate categories. Several instances from the Connections game require knowledge of morphology, phonology, or orthography for correct categorization. For example, as shown in Table 1, one needs knowledge about morphology to group Dom, Ion, Ness, and Ship as Noun Suffixes. Similarly, one needs to rely on phonological knowledge about the sound patterns of Answer, Two, Wrist, and Wrong to categorize them as Silent "W".
3.2.6 Combined
@@
Some of the hardest examples in the NYT Connections game require reasoning with multiple types of knowledge. For instance, the example in Table 1 shows that to group Deli, Niece, Roam, and Soul, one requires the knowledge that these words have the same phonological form with the cities Delhi, Nice, Rome, and Seoul. This categorization requires the simultaneous use of Encyclopedic and Linguistic Knowledge. Similarly, to group the words Book, Gram, In, and Tube together one needs to identify that they are essentially parts of closed compounds (Face+Book, Insta+Gram, Linked+In, You+Tube) that also represent popular social media apps. This categorization requires one to use Encyclopedic and Linguistic Knowledge together.

### v6

@@ Abstract
The New York Times Connections game has emerged as a popular and challenging pursuit for word puzzle enthusiasts. We collect 438 Connections games to evaluate the performance of state-of-the-art large language models (LLMs) against expert and novice human players. Our results show that even the best-performing LLM, Claude 3.5 Sonnet, which has otherwise shown impressive reasoning abilities on a wide variety of benchmarks, can only fully solve 18% of the games. Novice and expert players perform better than Claude 3.5 Sonnet, with expert human players significantly outperforming it. We create a taxonomy of the knowledge types required to successfully cluster and categorize words in the Connections game. We find that while LLMs perform relatively well on categorizing words based on semantic relations they struggle with other types of knowledge such as Encyclopedic Knowledge, Multiword Expressions or knowledge that combines both Word Form and Meaning. Our results establish the New York Times Connections game as a challenging benchmark for evaluating abstract reasoning capabilities in AI systems.
@@
Prisha Samadarshi1 Mariam Mustafa1 Anushka Kulkarni1 Raven Rothkopf1 Tuhin Chakrabarty2††thanks: Equal contribution. †denotes work done at Columbia† Smaranda Muresan1* 1 Department of Computer Science; Barnard College, Columbia University 2 Department of Computer Science, Stony Brook University {ps3203, mm5970, ajk2256, rgr2124}@alum.barnard.edu, tuhin.chakrabarty@stonybrook.edu, smuresan@barnard.edu
@@ 1 Introduction
Abstract reasoning represents a person’s ability to solve problems, identify patterns, and work with logical systems (Barrett et al., 2018; Johnson et al., 2021; Ji et al., 2022). While the performance of large language models (LLMs) on arithmetic and language-based commonsense reasoning benchmarks has been the subject of recent analyses, it is unclear whether these LLMs possess abstract reasoning capabilities that are often challenging even for humans (Xu et al., 2023). We propose the NYT Connections Game as a test bed for investigating the abstract reasoning capabilities of both humans and large language models (LLMs).
Connections is an engaging game launched by the New York Times (NYT) in June 2023. This daily game presents players with a 4x4 grid containing 16 words and tasks them with identifying four distinct clusters that link the corresponding four words in each cluster through some shared characteristics (Figures 1 [a] and [b]). Categories 1 (yellow), 2 (green), 3 (blue), and 4 (purple) are arranged in ascending level of difficulty. Category 1 is the most intuitive, while Category 4 is the hardest. For instance, in Figure 1 (b), the most straightforward category is "Conformists" {Followers, Lemmings, Puppets, Sheep}, while the most challenging category includes {Apartment, Insults, Likes, Shovels} and requires the understanding that a single word (in this case, "digs") can have multiple meanings that differ in etymology or sense, depending on the context.
While the task might seem easy, many words can be grouped easily into multiple categories, acting as red herrings. For instance, from the game in Figure 1, Likes, Followers, Shares, Insult might be categorized as “Social Media Interactions" at first glance. Unlike common categories (e.g., “Fruit," “Furniture”), the game is designed to promote ad hoc category formations that violate the correlational structure of the environment and are not well established in memory (Barsalou, 1983). To group words across proper categories, as shown in Figure 1 (b), a player must reason with various forms of knowledge spanning from Semantic Knowledge (Conformists) to Encyclopedic Knowledge (U.S. cities).
@@
We test the capabilities of five state-of-the-art large language models, namely Google’s Gemini 1.5 Pro (Team et al., 2023), Anthropic’s Claude 3.5 Sonnet (Anthropic, 2024), OpenAI’s GPT-4Omni (OpenAI, 2023), Meta’s Llama 3.1 405B, (AI@Meta, 2024) and Mistral Large 2 (Mistral-AI, 2024) on 438 distinct NYT Connections games and compare them with human performance on a subset of these games. Our experimental results show that while all LLMs can partially solve some games, their performance is far from ideal. Even the best-performing model Claude 3.5 Sonnet (with few-shot and chain-of-thought prompting), can only solve 18% of the games completely. In addition, we recruit human players at novice and expert levels of proficiency and compare their performance to Claude 3.5 Sonnet. Our results show that the NYT Connections game serves as a challenging benchmark for reasoning, with novice players performing marginally better than Claude 3.5 Sonnet and expert players performing significantly better than Claude 3.5 Sonnet in solving games perfectly (Section 5).
In addition, to better understand the LLMs abstract reasoning capabilities or the lack thereof, we propose a taxonomy of knowledge required to group words into their respective categories (Section 3.2). Our analysis at Section 6.1 shows that while LLMs are relatively better at reasoning that involve Semantic Relations, they struggle with other types of knowledge such as Multiword Expressions and combined knowledge about Word Form and Word Meaning (Section).
Our code and data will be made available to the public at https://github.com/mustafamariam/LLM-Connections-Solver.
@@ 2 Related Work
The growing popularity of Large Language Models has led to an exciting array of research using natural language processing techniques for text-based games. Recent work has studied whether these models can act as players in agentic environments Huang et al. (2024); Wang et al. (2023); Wu et al. (2023); Noever et al. (2020) or conversational settings Qiao et al. (2023). In addition to acting as players, researchers have also tested the abilities of transformer-based language models in generating games Ammanabrolu and Riedl (2021); Todd et al. (2023); Sudhakaran et al. (2024); Hu et al. (2024); Chen et al. (2023); Merino et al. (2024)
@@
Recent research has explored applying large language models (LLMs) and other natural language processing (NLP) techniques to solve and generate text-based puzzles. Wallace et al. (2022) propose automatic ways of solving crossword puzzles by generating answer candidates for each crossword clue using neural question answering models and combining loopy belief propagation with local search to find full puzzle solutions. Zhao and Anderson (2023) release PUZZLEQA, a multiple-choice dataset comprising 15 years of on-air Sunday Puzzle word-games and show that while ChatGPT can solve these questions with an accuracy of around 50%, they still struggle with generating novel and engaging puzzles. Unlike the NYT Connections game, PUZZLEQA relies on character-level word transformations compared to encyclopedic, associative, or semantic knowledge. Rozner et al. (2021) examined the potential of using "cryptic crossword" clues as an NLP benchmark.
@@
Most relevant to our paper is the contemporaneous work by Todd et al. (2024) who test the performance of various LLMs (from BERT and RoBERTA to GPT-4 and GPT-4 with Chain-of-Thought prompting) in solving the NYT Connections game. Our work is similar in that it utilizes the NYT Connections puzzles as a means to investigate the abstract reasoning capabilities of state-of-the-art LLMs. However, our contribution focuses not only on the ability of LLMs to solve the game, but also on studying the types of knowledge required to so. Moreover, we evaluate both state–of-the-art models (open and closed weights) and humans (novices and experts). We also benchmark on a larger number of NYT Connections game. Additionally, Todd et al. (2024)’s experimental setup mirrors the way the original NYT Connections game is played (i.e., one category at a time, with an allotment of 4 incorrect guesses), while we require both humans and LLMs to provide all categories at once in only one attempt.
@@
The word association task Galton (1879) has been used extensively in psychological and linguistic research as a way of measuring connections between words in the mental lexicon. Responses in word association tasks have informed what we know about the structure and organization of semantic memory and the mental lexicon De Deyne and Storms (2008). In this work, we similarly show how one must utilize semantic and associative memories to solve the NYT Connections game.
@@
Chollet (2019) proposed the Abstraction and Reasoning Corpus (ARC), built upon an explicit set of priors designed to be as close as possible to innate human priors and argued that it can be used to measure a human-like form of general fluid intelligence, enabling fair general intelligence comparisons between AI systems and humans. Recently Xu et al. (2023) show that GPT-4 solves only 13/50 of the most straightforward ARC tasks, demonstrating a significant gap in the abstract reasoning capabilities of LLMs. Prior work has also studied abstract reasoning in Neural Networks Barrett et al. (2018) in the presence of distracting features Zheng et al. (2019). Our work builds upon these and presents the NYT Connections game as a compelling benchmark for abstract reasoning capabilities of LLMs in the presence of distractors.
@@ 3.1 Collection
To gather the necessary data, we found an archival site consisting of all possible answer choices and their corresponding categorizations. As the NYT does not maintain an archive of NYT Connections puzzles, we resorted to an external, third-party site for data collection.111https://tryhardguides.com/nyt-connections-answers/ Our data spans daily problems from the conception of NYT Connections in June 2023 to August 2024. In total, we gather 441 distinct games, out of which 3 are used for few-shot prompting, while the remaining 438 comprise the dedicated test set.
@@ 3.2 Types of Reasoning
Investigating the relationship between words offers insights into both the structure of language and the influence of cognition on linguistic tasks (Stella et al., 2018). To solve the NYT Connections game, players must draw on certain aspects of word knowledge, such as a word’s meaning or form and sometimes both simultaneously. To deepen our understanding, we bucket each category, grouping into the types of knowledge that are primarily required to solve them. Three linguists annotate a total of 1,752 samples coming from 438 games into 3 broader categories which give rise to 8 subcategories (Figure 2). We take majority voting to arrive at a unique label for each category, grouping. We restrict annotations to sub-categories and not sub-sub-categories (e.g., Types of Semantic Relations). We obtain a Fleiss Kappa of 0.78 showing substantial agreement.
3.2.1 Word Form
Word Form refers to the specific shape or appearance a word takes in a given context. It encompasses various aspects of a word’s structure and representation and is broadly used in the NYT Connections game, testing knowledge on Phonology, Orthography, Morphology and Multiword Expressions. Phonology in Word Form deals with the sound structure of words, including pronunciation, stress patterns, and phonetic variations. Orthography relates to the conventional spelling system of a language, which may not always directly correspond to pronunciation. Morphology examines the internal structure of words, including roots and affixes and how they combine to create meaning. For example, as shown in Figure 2, one needs morphological knowledge to group Dom, Ion, Ness, and Ship as “Noun Suffixes." Similarly, one needs to rely on phonological knowledge of the sound patterns of Answer, Two, Wrist, and Wrong to categorize them as “Silent ‘W’." Multiword Expressions (MWE) have a fixed or semi-fixed form and are typically non-compositional (i.e., their meaning cannot be predicted from their individual components) Moon (1998).
3.2.2 Word Meaning
Semantic Relations:
@@ 3.2.2 Word Meaning
The majority of instances in the NYT Connections game require possessing knowledge of semantic relations Murphy (2003); Cruse (1986), such as synonymy (words with the same meaning), hypernymy/hyponymy (relation between a generic term and its specific instance), and homonymy and polysemy (many possible meanings of a word). Figure 2 shows three examples of groups that use such semantic relations.
Associative Relations
Prior work has studied models of automatic priming for word identification that are typically divided into groups based on associative relations (e.g., spreading activation) and others based on semantic similarity (e.g., distributed models) Thompson-Schill et al. (1998). In contrast to category members that share semantic features and category nodes, Associative Relations are elements of specific situations or thematic contexts, with little or no overlap between their semantic features (e.g. “Things that are red": Mars and Strawberry; Figure 2) Rose and Rahman (2016); Shanks (1995); Barsalou (1983).
Encyclopedic
@@
We notice that to group certain sets of words into their proper categories, one needs knowledge that spans beyond concepts and relies on entities in the real world found in knowledge bases such as Wikipedia Mihalcea and Csomai (2007). This can be seen in Figure 2, where, to bucket the words Globe, Mirror, Post, and Sun into the category of “Newspaper Names," one needs to possess knowledge that Globe refers to the Boston Globe, Mirror to the Daily Mirror, a UK tabloid, Post to the Washington Post and Sun to The Sun, another UK tabloid. We label this type of knowledge Encyclopedic.
3.2.3 Word Form + Word Meaning
@@
Some of the hardest examples in the NYT Connections game require reasoning of both word form and meaning. For instance, the example in Figure 2 shows that to group the words Book, Gram, In, and Tube, one needs to identify that they are essentially parts of closed compounds (Face+Book, Insta+Gram, Linked+In, You+Tube) that also represent popular social media apps. This categorization requires the use of knowledge on Word Meaning (Encyclopedic) as well as Word Form (Morphology).
| Word Form | Word Meaning |
Encyclopedic | 92 | ||||||||
| 44 | 168 | 1045 | 137 | 266 | |||||||||